# REST API Testing in DevOps: A Study on an Evolving Healthcare IoT Application

### v1

@@ 4.1 REST APIs Under Test
Table 1 presents the characteristics of all accessible APIs [^1] APIs for which the access provided by Oslo City as a part of experimental apparatus for research purposes. Inaccessible APIs are for sensitive data with privacy concerns. used in our experiments. Each API is associated with a specific functionality within the healthcare IoT application under test. These APIs consist of multiple endpoints, each targeting different types of features. The Alerts API has five features for assigning alerts, closing alerts, all alerts, total number of alerts, and dashboard alerts. The Authentication API targets different methods for authenticating different users. This includes authentication through email, password, SMS, phone number, device’s IMEI, token, medical device, and login credentials. In addition, it includes features such as password recovery for forgotten passwords, updating login credentials, and creating an emergency profile for a user. Three device APIs are related to three medicine dispensers (i.e., Karie [49], Medido [50], and Pilly [51]) that are integrated with the healthcare IoT application. Karie API provides features for modifying device settings, creating medication plans, and loading medical plans from pharmacies. Medido API enables modifying device settings and creating mock and real medication plans. Pilly API only allows for the configuration of device settings.
@@
The API for Patients targets features are related to patients’ health information, billings, time logs, health tracking, patients’ actions to perform, and reimbursements. The Users API focuses on user actions, that is, adding notes, creating events, threshold, searching users based on name or birth date, validating a user, updating user information, counting or listing down users, logging user interactions, and matching with existing patients. The Courses API is concerned with e-learning courses with features for course category, courses list, course modules, linking a course, uploading a new course, and finishing a course. The API for Mobile App targets features related to app version, app information, patient’s timeline, password information, and checking old and new passwords. The Reimbursements API includes features for creating billing summaries, reimbursements time logs, patient reimbursements status, measurements related reimbursements, task identifiers, and tracking reimbursements.
@@
The API for User Tasks provides features for various user tasks, i.e. assigning tasks to the concerned users, patient health surveys, task lists, task reminders, general tasks for administrative purposes, and global surveys for different users. The Invoicing API provides features for adding, canceling, replacing, terminating, and subscribing invoices. The Catalog API supplies catalogs with various types, models, units, features, and lists of catalogs. The API for Reports provides reports in different formats, patients’ reports, patients’ medication reports, zone-wise reports, and event-specific reports. The Care API is related to patient care messages, events, and reminder status. The Development API is concerned with creating smart alert rules, user profiles, and temporary interface generation.
@@
Table 1: Characteristics of APIs selected for the experiments, i.e., API label, number of endpoints (EPs), and features
API Label EPs Features
Alerts 5 Assign, Close, All, Total, Dashboard
@@
Authentication 13 Email, Password, SMS, Phone, IMEI, Token, Device, Log-in, Forgot password (Send/Resend), Credentials, Emergency profile
Device-Karie 3 Settings, Plan, Pharmacy
Device-Medido 3 Settings, Plan, Mock Plan
Device-Pilly 1 Settings
@@
Patients 6 Health data, Billings, Time logs, Tracking, Actions, Reimbursements
@@
Measurements 5 Single & Multiple inputs (Manual/Auto), Time series
@@
Users 14 Notes, Events, Threshold, Search methods, Validate, Update, Count/List, Logs, Match patient
@@
Courses 6 Category, List, Module, Link, Upload, Finish
@@
Mobile App 6 Version, Information, Timeline, Password info, Check old & new password
@@
Reimbursements 11 Billing summary, Time logs, Patients, Measurements, Task identifiers, Track
@@
User Tasks 9 Assignments, Survey, Tasks, Reminder, General tasks, Global survey
@@
Invoicing 9 Add, Cancel, Replace, Terminate, Subscriptions (Check, Fix, Aggregate)
Catalog 8 Types, Models, Units, Features, List
@@
Reports 10 Formats (HTML, JSON, PDF), Patients, Medication, Zone, Events
Care 4 Messages, Events, Reminder status
Development 7 Smart alert rules, Profiles, Auto UI generation
Total (APIs: 17) 120 -
@@ 4.2 Testing Tools Selection
Table 2 presents information for REST API testing tools selected for our experiments, including details such as tool name, version, release date, and applied testing techniques. The most recent version of each tool was picked as of May 10, 2024. To select tools for our experiments, we first analyzed empirical studies focusing on tool comparison. Drawing from the recommendations of effective and widely-used tools in previous studies [26, 48, 9], we initially shortlisted eight tools for our experiments, including RESTest, EvoMaster, RESTler, RESTCT, RestTestGen, bBOXRT, Morest, and Schemathesis. We later expanded our list to include TnT-Fuzzer, Got-Swag, RapiTest, QuickREST, and three recent tools: NLPtoREST, ARAT-RL, and foREST. Initially, we configured all tools to test the REST APIs of our application under test. Upon running these tools, we encountered various compatibility issues and crashes. After diagnosing these problems, we identified their common causes. Based on our investigation and insights from previous empirical studies [48, 9], we formulated a set of criteria required for a tool to qualify for testing our application under test. Note that resolving issues related to the tools could potentially introduce an internal validity threat, which also falls outside the scope of our paper. Next, we outline the criteria and elaborate on our systematic approach to the tool selection process.
@@
    Research-based tools derived from various research works with comprehensive and well-founded techniques (similar to earlier works [48, 9]).
@@
Among all available REST API testing tools, RESTCT only supports the OpenAPI specification version 2.0 (violating C1). Tools like TnT-Fuzzer and Got-Swag, as referenced in [5], also support OpenAPI version 2.0, thereby violating C1. The public version of Morest only provides a replication package, and its repository was last updated on September 13, 2021 (violating C2). bBOXRT provides an initial version as a replication package, last updated on January 30, 2020 (violating C2). In addition, bBOXRT requires an input API file in Java, which is not available in our case. In a similar line of work, RapiTest and QuickREST only provide replication packages, violating C2. Recent tools NLPtoREST and ARAT-RL provide replication packages, in the initial development stages, and can be run on Ubuntu (violating C2 & C3). The tool link for another recent black-box testing tool foREST [52] is also unavailable. Moreover, foREST is in early development phases, violating C2. Tools like Tcases, APIFuzzer, and Dredd are open-source tools without a research background, violating criterion C4. After careful inspection, we selected RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen for our experiments. Below is a brief overview of these selected REST API testing tools.
@@
RESTest [3] is an automated tool designed to test REST APIs of web applications, which work mainly in a black-box testing mode. This tool utilizes the API specification and test configurations as input to generate test cases. For test generation, RESTest supports various testing techniques, including constraint-based testing (CBT), adaptive random testing (ART), random testing (RT), and fuzz testing (FT). In addition, RESTest offers advanced features such as online testing in a production environment, handling inter-parameter dependencies, and generating realistic test data.
@@
EvoMaster [8] is a Web API testing tool that employs evolutionary techniques to generate test cases. It supports black-box and white-box testing modes and is compatible with various API types, including REST and GraphQL. In black-box testing, EvoMaster takes the API specification as input. White-box testing requires additional inputs, including source code and a driver. For black-box mode, EvoMaster employs fuzzing techniques for test generation, while in white-box mode, it utilizes many objective search algorithms for test generation.
@@
Schemathesis [5] is a web API fuzz testing tool that utilizes property-based testing for generating test cases. It operates in black-box and white-box testing modes and is compatible with APIs specified in OpenAPI or GraphQL formats. This tool is designed to test cross-platform web APIs, irrespective of the programming language used for development. In addition, Schemathesis has been used in various industrial projects, thereby streamlining the process of automated REST API testing.
@@
RESTler [4] is a web services fuzz testing tool mainly focusing on identifying security and reliability faults in REST APIs. It takes API specification in OpenAPI format, analyzes the specification to identify dependencies among requests, generates fuzz tests, and executes tests on the API to explore web services. Based on responses originating in various test executions, it learns distinct bug patterns. It supports two types of fuzzing, i.e., fuzz and fuzz-lean.
@@
RestTestGen [7] is a REST API testing tool with a primary focus on black-box testing. It supports two testing techniques: nominal and error testing and mass assignment security testing. RestTestGen offers an extensible interface that facilitates adding new features and implementing novel testing techniques. In this context, a recent tool NLP2REST [34] is integrated with RestTestGen, leveraging natural language processing techniques to enhance functionality in API testing scenarios.
@@ 4.3.1 Experiment Preparation
Our experiments involve five REST API testing tools, each supporting one or more testing techniques. For each testing mode, we set up different configurations of these tools. RESTest configured with the CBT, ART, RT and FT testing techniques is termed RESTest-CBT, RESTest-ART, RESTest-RT and RESTest-FT, abbreviated as RT-CBT, RT-ART, RT-RT and RT-FT, respectively. EvoMaster configured with back-box testing mode is termed EvoMaster-BB (in short EM-BB). Schemathesis configured with property-based testing mode is termed Schemathesis-PT (in short ST-PT). RESTler configured with fuzz and fuzz-lean testing modes is termed RESTler-F and RESTler-FL, abbreviated as RL-Fz and RL-FL, respectively. RestTestGen configured with nominal and error testing and mass assignment security testing is termed RestTestGen-NET and RestTestGen-MAST, abbreviated as RTG-NET and RTG-MAST respectively. These configurations result in a total of 10 tools for our experiments.
@@
Figure 3 shows the design of the experiment. We started with the first weekly release (R1) and ran all testing tools on that release. At the end of execution, we collected test cases and results produced by each testing tool for R1. For each next release, R2-R14, we executed the same set of tests generated by each tool on each subsequent release. One of the primary objectives of our experiment is to identify regressions. To achieve this, using the same set of tests (generated from each tool) is necessary. This is due to the inherent randomness factor involved in each testing tool, leading to potential variability in the tests generated at different times. For example, the tests generated during the initial run may differ from those generated during subsequent runs. This variability could hinder the detection of regressions in evolving releases. Therefore, to effectively identify regressions across different releases, we used the set of tests generated for R1 to test releases R2-R14. This approach to regression testing is commonly referred to as the retest-all strategy in literature [53, 54]. We chose this strategy because it does not necessitate a test selection, minimization, or prioritization approach to determine which specific test cases should be rerun. To the best of our knowledge, no such approach exists for REST API test cases. Furthermore, adapting existing test selection, minimization, or prioritization approaches for test cases created for REST APIs could potentially introduce significant bias, such as implementation bias [16]. This also deviates from the primary focus of our study.
@@ 4.3.2 Experiment Execution
To account for the inherent randomness of REST API testing tools, we executed each tool 10 times in parallel, with each run limited to one hour. Furthermore, our experiments involve three real medical devices in the loop. Given the constraints on the ability to handle requests from these devices and the potential damage to the device [1, 15], we introduced a delay of three seconds for APIs that require interaction with medical devices. This execution setup aligns with that used in our previous work [13].
@@
Table 3: Evaluation aspect and analyses metrics, measures, and statistical tests corresponding to each RQ
RQ Aspect Analysis Metrics, Measures, and Statistical Tests
RQ1 Failure Gestalt similarity, Unique failures, Wilcoxon test, Vargha-Delaney Â12 measure
RQ2 Fault Gestalt similarity, Unique faults
RQ3 Coverage Percentage coverage, Wilcoxon test, Vargha-Delaney Â12 measure

### v2

@@ 4.1 REST APIs Under Test
Table 1 presents the characteristics of all accessible APIs [^1] APIs for which the access was provided by Oslo City as a part of an experimental apparatus for research purposes. Inaccessible APIs are for sensitive data with privacy concerns. used in our experiments. These APIs are part of the staging environment set up by Oslo City’s industry partner specifically for testing purposes. Each API is associated with a specific functionality within the healthcare IoT application under test. These APIs consist of multiple endpoints, each utilizing different HTTP request methods and targeting various features. As observed in Table 1, most endpoints primarily use GET and POST methods for CRUD operations. However, requests such as PUT and DELETE, which involve creating and deleting resources, are inaccessible to us to ensure controlled access for testing and prevent unintended changes to the application’s functionality. The Alerts API has five features for assigning alerts, closing alerts, all alerts, the total number of alerts, and dashboard alerts. The Authentication API targets different methods for authenticating different users. This includes authentication through email, password, SMS, phone number, device’s IMEI, token, medical device, and login credentials. In addition, it includes features such as password recovery for forgotten passwords, updating login credentials, and creating an emergency profile for a user. Three device APIs are related to three medicine dispensers (i.e., Karie karie , Medido medido , and Pilly pilly ) that are integrated with the healthcare IoT application. Karie API provides features for modifying device settings, creating medication plans, and loading medical plans from pharmacies. Medido API enables modifying device settings and creating mock and real medication plans. Pilly API only allows for the configuration of device settings.
@@
The API for Patients targets features are related to patients’ health information, billings, time logs, health tracking, patients’ actions to perform, and reimbursements. The Users API focuses on user actions, that is, adding notes, creating events, threshold, searching users based on name or birth date, validating a user, updating user information, counting or listing down users, logging user interactions, and matching with existing patients. The Courses API is concerned with e-learning courses with features for course category, courses list, course modules, linking a course, uploading a new course, and finishing a course. The API for Mobile App targets features related to app version, app information, patient’s timeline, password information, and checking old and new passwords. The Reimbursements API includes features for creating billing summaries, reimbursement time logs, patient reimbursement status, measurements related reimbursements, task identifiers, and tracking reimbursements.
@@
The API for User Tasks provides features for various user tasks, i.e., assigning tasks to the concerned users, patient health surveys, task lists, task reminders, general tasks for administrative purposes, and global surveys for different users. The Invoicing API provides features for adding, canceling, replacing, terminating, and subscribing to invoices. The Catalog API supplies catalogs with various types, models, units, features, and lists of catalogs. The API for Reports provides reports in different formats, patients’ reports, patients’ medication reports, zone-wise reports, and event-specific reports. The Care API is related to patient care messages, events, and reminder status. The Development API is concerned with creating smart alert rules, user profiles, and temporary interface generation.
@@
Table 1: Characteristics of APIs selected for the experiments, i.e., API label, number of endpoints (EPs), EPs corresponding to HTTP request methods, and features
API Label EPs EPs w.r.t. Requests Features
Alerts 5 2 POST, 3 GET Assign, Close, All, Total, Dashboard
@@
Authentication 13 13 POST Email, Password, SMS, Phone, IMEI, Token, Device, Log-in, Forgot password (Send/Resend), Credentials, Emergency profile
Device-Karie 3 1 GET+POST, 2 POST Settings, Plan, Pharmacy
Device-Medido 3 2 GET+POST, 1 GET Settings, Plan, Mock Plan
Device-Pilly 1 1 GET+POST Settings
@@
Patients 6 6 POST Health data, Billings, Time logs, Tracking, Actions, Reimbursements
@@
Measurements 5 5 POST Single & Multiple inputs (Manual/Auto), Time series
@@
Users 14 1 GET, 13 POST, 1 GET+POST Notes, Events, Threshold, Search methods, Validate, Update, Count/List, Logs, Match patient
@@
Courses 6 2 GET, 1 POST, 3 GET+POST Category, List, Module, Link, Upload, Finish
@@
Mobile App 6 2 POST, 4 GET Version, Information, Timeline, Password info, Check old & new password
@@
Reimbursements 11 3 POST, 8 GET Billing summary, Time logs, Patients, Measurements, Task identifiers, Track
@@
User Tasks 9 9 GET+POST Assignments, Survey, Tasks, Reminder, General tasks, Global survey
@@
Invoicing 9 5 POST, 4 GET Add, Cancel, Replace, Terminate, Subscriptions (Check, Fix, Aggregate)
Catalog 8 8 GET Types, Models, Units, Features, List
@@
Reports 10 1 POST, 9 GET Formats (HTML, JSON, PDF), Patients, Medication, Zone, Events
Care 4 4 GET Messages, Events, Reminder status
Development 7 7 GET Smart alert rules, Profiles, Auto UI generation
Total (APIs: 17) 120 - -
@@ 4.2 Testing Tools Selection
Table 2 presents information for REST API testing tools selected for our experiments, including details such as tool name, version, release date, and applied testing techniques. The most recent version of each tool was picked as of May 10, 2024. To select tools for our experiments, we first analyzed empirical studies focusing on tool comparison. Drawing from the recommendations of effective and widely-used tools in previous studies golmohammadi2022testing ; zhang2023open ; kim2022automated , we initially shortlisted eight tools for our experiments, including RESTest, EvoMaster, RESTler, RESTCT, RestTestGen, bBOXRT, Morest, and Schemathesis. We later expanded our list to include TnT-Fuzzer, Got-Swag, RapiTest, QuickREST, and three recent tools: NLPtoREST, ARAT-RL, and foREST. Initially, we configured all tools to test the REST APIs of our application under test. Upon running these tools, we encountered various compatibility issues and crashes. After diagnosing these problems, we identified their common causes. Based on our investigation and insights from previous empirical studies zhang2023open ; kim2022automated , we formulated a set of criteria required for a tool to qualify for testing our application under test. Note that resolving issues related to the tools could potentially introduce an internal validity threat, which also falls outside the scope of our paper. Next, we outline the criteria and elaborate on our systematic approach to the tool selection process.
@@
    Research-based tools derived from various research works with comprehensive and well-founded techniques (similar to earlier works zhang2023open ; kim2022automated ).
@@
Among all available REST API testing tools, RESTCT only supports the OpenAPI specification version 2.0 (violating C1). Tools like TnT-Fuzzer and Got-Swag, as referenced in hatfield2022deriving , also support OpenAPI version 2.0, thereby violating C1. The public version of Morest only provides a replication package, and its repository was last updated on September 13, 2021 (violating C2). bBOXRT provides an initial version as a replication package, last updated on January 30, 2020 (violating C2). In addition, bBOXRT requires an input API file in Java, which is not available in our case. In a similar line of work, RapiTest and QuickREST only provide replication packages, violating C2. Recent tools NLPtoREST kim2023enhancing , ARAT-RL kim2023adaptive , APIRL foley2025apirl , and LlamaRestTest kim2025llamaresttest provide replication packages, are still in initial development stages, with NLPtoREST, ARAT-RL, and LlamaRestTest necessitating execution on Ubuntu (violating C2 & C3). The tool link for another recent black-box testing tool, foREST lin2023forest , is also unavailable. Moreover, foREST is in early development phases, violating C2. Tools like Tcases, APIFuzzer, and Dredd are open-source tools without a research background, violating criterion C4. After careful inspection, we selected RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen for our experiments. Below is a brief overview of these selected REST API testing tools.
@@
RESTest martin2021restest is an automated tool designed to test REST APIs of web applications, which work mainly in a black-box testing mode. This tool utilizes the API specification and test configurations as input to generate test cases. For test generation, RESTest supports various testing techniques, including constraint-based testing (CBT), adaptive random testing (ART), random testing (RT), and fuzz testing (FT). In addition, RESTest offers advanced features such as online testing in a production environment, handling inter-parameter dependencies, and generating realistic test data. These features are enabled through various techniques such as the CBT approach for handling inter-parameter dependencies martin2020restest ; martin2021specification .
@@
EvoMaster arcuri2018evomaster is a Web API testing tool that employs evolutionary techniques to generate test cases. It supports black-box and white-box testing modes and is compatible with various API types, including REST and GraphQL. In black-box testing, EvoMaster takes the API specification as input. White-box testing requires additional inputs, including source code and a driver. For black-box mode, EvoMaster employs fuzzing techniques for test generation, while in white-box mode, it utilizes many objective search algorithms for test generation.
@@
Schemathesis hatfield2022deriving is a web API fuzz testing tool that utilizes property-based testing for generating test cases. It operates in black-box and white-box testing modes and is compatible with APIs specified in OpenAPI or GraphQL formats. This tool is designed to test cross-platform web APIs, irrespective of the programming language used for development. In addition, Schemathesis has been used in various industrial projects, thereby streamlining the process of automated REST API testing.
@@
RESTler atlidakis2019restler is a web services fuzz testing tool mainly focusing on identifying security and reliability faults in REST APIs. It takes an API specification in OpenAPI format, analyzes the specification to identify dependencies among requests, generates fuzz tests, and executes tests on the API to explore web services. Based on responses originating from various test executions, it learns distinct bug patterns. It supports two types of fuzzing, i.e., fuzz and fuzz-lean.
@@
RestTestGen corradini2022resttestgen is a REST API testing tool with a primary focus on black-box testing. It supports two testing techniques: nominal and error testing, and mass assignment security testing. RestTestGen offers an extensible interface that facilitates adding new features and implementing novel testing techniques. In this context, a recent tool, NLP2REST kim2023enhancing , is integrated with RestTestGen, leveraging natural language processing techniques to enhance functionality in API testing scenarios.
@@ 4.3.1 Experiment Preparation
Our experiments involve five REST API testing tools, each supporting one or more testing techniques. For each testing mode, we set up different configurations of these tools. RESTest configured with the CBT, ART, RT, and FT testing techniques is termed RESTest-CBT, RESTest-ART, RESTest-RT, and RESTest-FT, abbreviated as RT-CBT, RT-ART, RT-RT, and RT-FT, respectively. EvoMaster configured with back-box testing mode is termed EvoMaster-BB (in short EM-BB). Schemathesis configured with property-based testing mode is termed Schemathesis-PT (in short ST-PT). RESTler configured with fuzz and fuzz-lean testing modes is termed RESTler-F and RESTler-FL, abbreviated as RL-Fz and RL-FL, respectively. RestTestGen configured with nominal and error testing and mass assignment security testing is termed RestTestGen-NET and RestTestGen-MAST, abbreviated as RTG-NET and RTG-MAST respectively. These configurations result in a total of 10 tools for our experiments.
@@
Figure 3 shows the design of the experiment. We started with the first weekly release (R1) and ran all testing tools on that release. For each tool’s execution on R1, we set up a proxy server to collect HTTP requests generated by each tool and responses obtained from the APIs. Note that instead of relying on third-party proxy solutions, which could introduce execution overhead or raise data privacy concerns for the industry partner, we developed a lightweight proxy to ensure minimal interference and precise data capture. To enable consistent test replay across subsequent releases, we also implemented a test runner that sequentially picks up each HTTP request initially generated by a tool, sends it to the appropriate API endpoint, and records the resulting response. Since the APIs under test only use GET and POST methods, as shown in Table 1 and discussed in Section 4.1, managing dynamic data dependencies like resource IDs—typically required for methods such as CREATE and DELETE—was not applicable in our study. Therefore, we did not implement adaptations to handle such dependencies in the replay test runner. However, in scenarios involving methods like CREATE and DELETE, a potential approach could be to maintain a dynamic map to track resource creation and deletion, ensuring consistent reuse of resource IDs across operations, regardless of the request execution order. This allowed us to reuse the same set of test cases across all subsequent releases (R2–R14), ensuring consistency in regression testing across evolving versions. Reusing the same test set is essential to the inherent randomness factor involved in each testing tool, leading to potential variability in the tests generated at different times. For example, the tests generated during the initial run may differ from those generated during subsequent runs. This variability could hinder the detection of regressions in evolving releases. Therefore, to effectively identify regressions across different releases, we used the set of tests generated for R1 to test releases R2–R14. This approach to regression testing is commonly referred to as the retest-all strategy in literature yoo2012regression ; greca2023state . We chose this strategy because it does not necessitate a test selection, minimization, or prioritization approach to determine which specific test cases should be rerun. To the best of our knowledge, no such approach exists for REST API test cases. Furthermore, adapting existing test selection, minimization, or prioritization approaches for test cases created for REST APIs could potentially introduce significant bias, such as implementation bias sartaj2024uncertainty . This also deviates from the primary focus of our study.
One of the potential implications of using the retest-all strategy in our experiments is simplicity in test generation and consistency in test execution across releases, which aligns with our experimental goals. However, other potential implications include increased resource consumption and redundancy due to rerunning all tests. It is important to note that for the APIs under test, we lacked prior knowledge of planned feature changes (e.g., additions, removals, or updates) across releases. Without such information and the availability of automated test selection, minimization, or prioritization approaches, we relied on the retest-all strategy for regression testing. This involved generating automated test cases once for R1 using each tool and reusing them for subsequent releases. If some tools supported these functions, our evaluation could have been designed accordingly. Instead of rerunning all tests from R1, we could have used test selection or minimization techniques to execute a targeted or reduced set of tests for each subsequent release, reducing resource consumption and avoiding redundancy. Another important consideration is the maintenance of the test suite, which requires regular review and evolution to align with API changes during DevOps. Although developing a dedicated approach for automated test suite maintenance is essential, it falls beyond the scope of our study and presents a promising direction for future research.
@@ 4.3.2 Experiment Execution
To account for the inherent randomness of REST API testing tools, we executed each tool 10 times in parallel, with each run limited to one hour. Our parallel execution setup was inspired by the existing work on tool comparison zhang2023open . Specifically, we configured eight threads to run eight parallel jobs using eight logical processors, ensuring that each run adhered to the one-hour limit and was forcibly terminated if it exceeded this duration. During execution, we observed that all tools consistently generated tests and utilized the full time budget, except RestTestGen-MAST, which did not generate any tests (as shown in Table 4) and terminated early without errors. Since no tests were produced (as detailed in Section 4.5), rerunning RestTestGen-MAST to fully utilize the allocated time—similar to the setup in zhang2023open —would have unnecessarily consumed resources. Therefore, the freed-up resources became available for other tools.
Running tools in parallel continuously can potentially lead to APIs becoming overloaded. To avoid this problem, we coordinated with Oslo City’s team and their industry partner while planning our experiments. They provided access to a dedicated staging environment, likely similar to the one used internally for testing. We communicated our testing schedule, estimated execution durations, and expected API call overload. During planned testing periods, they configured the API servers to efficiently manage the traffic generated from our designated test account to ensure smooth execution without affecting other system functionalities or users. Furthermore, our experiments involve three real medical devices in the loop. Given the constraints on the ability to handle requests from these devices and the potential damage to the device sartaj2023hita ; sartaj2024modelbased , we introduced a delay of three seconds for APIs that require interaction with medical devices. This execution setup for device APIs reduced API overload and aligns with the setup used in our previous work sartaj2023testing .
@@
Table 3: Evaluation aspect and analyses metrics, statistical tests, and measures corresponding to each RQ
RQ Aspect Analysis Metrics Statistical Tests Measures
RQ1 Failure Gestalt similarity, Total failures, Ratio Wilcoxon test Vargha-Delaney Â12
RQ2 Fault Gestalt similarity, Unique faults, Ratio, APFD
RQ3 Coverage Percentage coverage Wilcoxon test Vargha-Delaney Â12
@@
Figure 4: Example comparison of error responses in JSON format, highlighting one property with different values while other property values are the same.